One method of removing CO[sub]2[/sub] gas from a spacecraft is to allow the CO[sub]2[/sub] to react with LiOH.
How many liters of CO[sub]2[/sub] at 25.9 degrees Celcius and 751 Torr can be removed per kilogram of LiOH consumed?
1
Expert's answer
2010-11-22T16:46:49-0500
CO2 + 2LiOH = Li2CO3 + H2O
We have 1kg LiOH : 1000g/(7+16+1) g/mol = 1000/24 = 41.67 mol.


Thus M(CO2) = 2 M(LiOH) = 83.33 mol
m(CO2) = 83.33 mol * (12 + 2*16) g/mol = 4000 g = 4 kg.

Use the equation
pV = ν R T



p = 751 torr = 100 125.099 Pa, T = 25.9 C = 298.9 K
V = ν R T / p = 83.33 * 8.31 *298.9 /100125.099 = 2.07 m3 = 2067.2 Liters